# Thermal evolution of dark matter in the early universe from a symplectic glueball model

### v1

@@ 4 Interpretation as a model for dark matter
As the universe expands and cools down, the equilibrium thermodynamic quantities will decrease as discussed in section 3; when the temperature drops below $T_{\mbox{\tiny{c}}}$, the dark sector undergoes a qualitative change of state and the dark gluons confine into massive glueballs. This is accompanied by a brisk drop in the equilibrium thermodynamic quantities. In particular the change in the entropy density across $T_{\mbox{\tiny{c}}}$ can be interpreted in terms of the sharp reduction in the number of (effective) degrees of freedom of the theory: from $N(2N+1)=10$ “gluons”, with two polarizations each, in the high-temperature phase, to $O(1)$ massive “glueballs”. The finiteness of the latent heat implies that this confining transition is a discontinuous one.
In the low-temperature phase, the thermodynamics can be modelled in terms of a gas of massive glueballs: after confinement, a freeze-out process takes place, leading to a gas dominated by the contribution of the lightest state in the spectrum, i.e., the scalar glueball. If the dynamics of the latter can be described in terms of an effective Lagrangian of the form
which accounts for number-changing glueball self-interaction processes like $3\to 2$, then the number density of the glueball is expected to decrease as a function of time. While the $\lambda_{n}$ coefficients appearing in eq. (4.2) are a priori unknown, one can extract information about them by simple arguments. In particular, from naïve dimensional analysis follows that the energy dimension of the generic $\lambda_{n}$ coefficient is $4-n$. A slightly less trivial observation is that, since $\phi$ represents a glueball state in a Yang-Mills theory in which the “color” multiplicity of the dark gluons is $10$, one could invoke large-$N$ arguments [205] (see also ref. [206, section 2.2]) to argue that $\lambda_{n}=O(N^{2-n})$, so that the dominant glueball self-interaction processes are encoded in the cubic term appearing in eq. (4.2). As long as the self-interaction rate is larger than the expansion rate of the universe, the comoving entropy density of the gas of dark glueballs is conserved. Noting that the results for the glueball masses in table 1 and the value of the $\frac{T_{\mbox{\tiny{c}}}}{\sqrt{\sigma}}$ ratio reported in ref. [34] indicate that in the confining phase one can use a non-relativistic approximation for the equation of state, and including only the leading contribution for the lightest glueball (denoting its mass as $m$) the entropy density reduces to
$s\simeq m^{2}\sqrt{\frac{mT}{8\pi^{3}}}\exp\left(-\frac{m}{T}\right).$ (4.3)
As a consequence, by denoting the Robertson–Walker scale factor of the universe as $a$, entropy conservation in a comoving volume $a^{3}$ implies that the temperature of dark matter decreases only logarithmically with $a$:
$T\simeq\frac{m}{3\ln(a/a_{0})},$ (4.4)
The discontinuous nature of the confining transition implies that the energy-momentum tensor can experience large anisotropic fluctuations, possibly leading to the generation of gravitational waves, as was already suggested long ago [207, 208, 209, 210].
@@
The generation of gravitational waves can be described by the (linearized) Einstein equation for tensor perturbations in a Friedmann background; following the notations and conventions of ref. [211], and writing the metric as $ds^{2}=a^{2}(t)\left(-dt^{2}+d{\mathbf{x}}^{2}\right)$, the equation
@@
where dots denote derivatives with respect to the conformal time, $G$ is the gravitational constant, $\rho$ is the background cosmological energy density, $\Pi_{\pm}$ are the tensor helicity modes of the (dimensionless) anisotropic stress sourcing the gravitational waves, and $h_{\pm}$ represent the helicity modes of the gravitational waves. There are different mechanisms, that could be involved in the production of gravitational waves (see, for example, ref. [212, section 8] for a review): the simplest of these is the collision of true-vacuum bubbles [213, 214, 215]. Bubbles expanding in a thermal medium, however, are subject to friction: this slows them down to an approximately constant speed, and the energy transferred to the plasma generates sound waves, which can be another source of gravitational waves [209, 216, 217, 218, 219]. In addition, the bubble-merging process can also induce hydrodynamic turbulence, which is also a potential source of gravitational waves [220, 221, 222, 223, 224, 225].
After production, the propagation of gravitational waves is essentially free, with the phenomenologically very interesting implication that their experimental detection could provide a unique test of gravitational interactions of dark matter in the early universe. More precisely, the frequency of gravitational waves produced at the phase transition decreases as the inverse of the Robertson–Walker scale factor, while their fraction of the critical energy density decreases as $a^{-4}$. Following the discussion in ref. [37], one can then derive that entropy conservation implies that
@@
where $T_{\mbox{\tiny{CMB}}}=2.34864(49)\cdot 10^{-13}$ GeV denotes the temperature of the cosmic microwave background, while $g_{s}$ is the effective number of relativistic degrees of freedom, the $\star$ subscript refers to the time of production, while the 00 subscript refers to the time of observation.
The spectrum of gravitational waves produced by bubble collisions can be approximated as [215, 226, 211]
$\frac{d\Omega^{\mbox{\tiny{(B)}}}_{\rm GW}h^{2}}{d\ln k}\simeq\frac{2}{3\pi}h^{2}\Omega_{\mbox{\tiny{r}}0}{\cal H}_{*}^{2}t_{\mbox{\tiny{pt}}}^{2}\Omega_{S*}^{2}v^{3}\frac{(kt_{\mbox{\tiny{pt}}})^{3}}{1+(kt_{\mbox{\tiny{pt}}})^{4}},$ (4.7)
where $k=2\pi af$ and ${\cal H}=aH$ respectively denote the conformal wave number and the conformal Hubble parameter (i.e., the wave number and the Hubble parameter in units of $a^{-1}$), $\Omega_{\mbox{\tiny{r}}}$ is the radiation energy density, $t_{\mbox{\tiny{pt}}}$ denotes the duration of the phase transition, $v$ is the bubble velocity and $\Omega_{S*}=\rho_{S*}/\rho_{*,\rm crit}=\Omega_{*\rm GW}$ is the relative energy density in the source. Note that the duration of the phase transition can be defined in terms of the bubble nucleation rate $\Gamma$ via [227]
@@
and that the spectrum of gravitational waves produced by bubble collisions depends on the temperature of the phase transition through the dependence of ${\cal H}_{*}$ on $T_{*}$.
@@
As for the spectrum of gravitational waves produced by turbulence, one has [228, 37]
@@
$\frac{d\Omega^{\mbox{\tiny{(HD)}}}_{\rm GW}h^{2}}{d\ln k}$ $\simeq\frac{8}{\pi^{6}}h^{2}\Omega_{r0}{\cal H}_{*}t_{\mbox{\tiny{pt}}}\Omega_{S*}^{3/2}v^{4}\,\frac{(kt_{\mbox{\tiny{pt}}})^{3}}{(1+4k/{\cal H}_{*})\left[1+(vkt_{\mbox{\tiny{pt}}}/\pi^{2})\right]^{11/3}}.$ (4.9)
@@
A non-perturbative determination of quantities related to time-dependent phenomena and/or beyond thermodynamic equilibrium from lattice simulations is possible, but it requires techniques beyond those used for the equation of state. While Monte Carlo simulations of real-time dynamics would be hampered by a numerical sign problem, one can resort to an alternative method, focused on the study of a low-energy effective field theory describing the long-wavelength modes only [^6] A similar approach was proposed in ref. [229] and later implemented in refs. [230, 231, 232] to study the transverse-momentum broadening associated with the jet quenching phenomenon in the QCD plasma. through Langevin equations [233]: this approach was used in refs. [234, 227] to study the bubble nucleation rate (and the duration of the phase transition, which can be obtained from the latter via eq. (4.8)) at the electroweak phase transition. As discussed in ref. [37], the value of the gravitational-wave frequency corresponding to the maximum in signal intensity observable today depends on $t_{\mbox{\tiny{pt}}}$, as does the value of the signal-intensity maximum itself.

### v2

@@ 4 Interpretation as a model for dark matter and gravitational-wave production
Confining transition. As the universe expands and cools down, the equilibrium thermodynamic quantities will decrease as discussed in section 3; when its temperature drops below $T_{\mbox{\tiny{c}}}$, the dark sector undergoes a change of state, and the dark gluons confine into massive glueballs. This is accompanied by a brisk drop in the equilibrium thermodynamic quantities. In particular the change in the entropy density across $T_{\mbox{\tiny{c}}}$ can be interpreted in terms of the sharp reduction in the number of (effective) degrees of freedom of the theory: from $N(2N+1)=10$ “gluons”, with two polarizations each, in the high-temperature phase, to $O(1)$ massive “glueballs”. The finiteness of the latent heat implies that this confining transition is a discontinuous one.
Generation of gravitational waves at the phase transition. Gravitational-wave astronomy entered the domain of experimental science almost a decade ago [43]: since then, the Advanced Virgo [235] and the Advanced Laser Interferometer Gravitational-Wave Observatory [236, 237] have observed a large number of events consistent with gravitational waves produced by transient astrophysical sources: mergers of binary black holes or coalescence of binary neutron stars. A more ambitious goal for the future [238] consists in the detection of gravitational waves produced by cosmological sources [239, 240, 241]. In particular, novel ground-based interferometers of third generation, like the Einstein Telescope [242] and the Cosmic Explorer [243], or space-based gravitational wave detectors, like the Laser Interferometer Space Antenna (LISA) [49], the Big Bang Observer (BBO) [244], Taiji [48], TianQin [47], and DECIGO [245] could detect gravitational waves produced by violent phenomena in the early universe, among which strong cosmological phase transitions [246, 247, 40]. The discontinuous nature of a strong first-order phase transition, like the confining phase transition in our $\operatorname{\mathrm{Sp}}(2)$ model, implies that the energy-momentum tensor can experience large anisotropic fluctuations, resulting into the generation of a stochastic, “cosmological” gravitational-wave background: a phenomenon that was already suggested long ago [248, 249, 250, 251].
At this point, however, an important remark is in order: the estimates for the spectrum of gravitational waves that will be discussed in this section depend on parameters (including, in particular, the bubble wall velocity and the transition-strength parameter $\alpha$) for which we must necessarily rely on—sometimes relatively crude—estimates, that often involve underlying assumptions. This can be contrasted with the results for the equation of state and for the other quantities discussed in this work, for which we could derive robust ab initio predictions. Ultimately, the reason of this discrepancy lies in the very nature of the phenomena they are related to (and in the way they are studied in the numerical calculation). On the one hand, the pressure, the energy and entropy densities, etc. are equilibrium thermodynamic quantities, which have a natural formulation in terms of Euclidean quantum field theory, and for which a non-perturbative evaluation on a Euclidean lattice is in principle straightforward. On the other hand, quantities related to processes such as the formation, expansion, collision, etc. of bubbles of different phases within a medium encode information about phenomena beyond thermodynamic equilibrium and intrinsically dynamical. In general, the non-perturbative numerical evaluation of such quantities for a quantum field theory regularized on a lattice requires methods that are quite different from those that can be applied to study equilibrium thermodynamics; these include real-time simulations [252] or techniques to solve mathematically non-trivial inverse problems [253], which entail their own peculiar challenges. In cases when the quantum fluctuations of fields are small compared to the thermal ones, one can resort to classical statistical lattice gauge theory simulations, whereby the gauge fields are let evolve in a deterministic way, according to the classical equations of motion, and the process is repeated many times, starting from initial values for the fields sampled from a thermal distribution. This approach allows one to estimate the average value and the variance of various real-time quantities, and has been used in studies of non-equilibrium processes taking places in the early evolution of the system created immediately after a “small bang”, namely a heavy-ion collision [254].
A particularly important quantity associated with the bubble expansion is the bubble wall velocity, denoted by $v$, according to which one can distinguish among deflagrations (when the bubble wall velocity is smaller than the speed of sound, and a “sound shell” can form in front of the wall), detonations (when the bubble wall velocity is larger than the speed of sound, and the plasma behind the wall can have non-zero velocity), or hybrid cases. The bubble wall velocity is determined by the competing forces acting on the wall: the latent heat that is released in the phase transition pushes the bubble outwards, while its interaction with the plasma creates friction. When the two forces are balanced, the wall ceases to accelerate. Deriving an accurate estimate for the bubble wall velocity in a certain model, however, is highly non-trivial, as it cannot be determined only from the effective potential associated with a first-order transition, and requires instead knowledge of the friction of the plasma near the bubble wall, which can be determined only if the precise hydrodynamical solution of the plasma profile is accessible. While lattice studies of this problem do exist (see, for instance, ref. [255] as a recent example), they typically rely on a setup which is entirely different from the one of the lattice calculations reported in the present work—which, instead, are based on the “microscopic” formulation of the theory in terms of fundamental fields. Extending our work to include this kind of lattice calculations lies beyond the scope of the present study and would entail significant computational costs, hence we leave it for the future. It is nevertheless possible to provide at least a rough, order-of-magnitude estimate of $v$, based on the following argument [248, 256, 257]: let $x$ denote the fraction of the dark sector in the confining phase. By the definition of the latent heat, when this fraction increases by an amount $dx$, this is accompanied by the release (the confining transition being an exothermic process) of an amount of energy per volume $L_{\mbox{\tiny{h}}}dx$, which corresponds to a temperature variation
$dT=\frac{L_{\mbox{\tiny{h}}}}{\frac{d\epsilon}{dT}}dx=\frac{15}{2\pi^{2}g^{\mbox{\tiny{eff}}}_{\star}}\frac{L_{\mbox{\tiny{h}}}}{T^{3}}dx,$ (4.2)
where $g^{\mbox{\tiny{eff}}}_{\star}$ denotes the effective number of relativistic degrees of freedom in the thermal bath: for the dark sector, it can be expressed as the product of the number of gluon polarizations ($2$) times the multiplicity factor associated with the dimension of the Lie algebra of the gauge group (which is $10$ for the $\operatorname{\mathrm{Sp}}(2)$ theory, or $N(2N+1)$ for a generic $\operatorname{\mathrm{Sp}}(N)$ theory), times a coefficient of order 11—encoding the fact that, as our results for the equation of state show, at temperatures of the order of $T_{\mbox{\tiny{c}}}$ the energy density in the deconfined phase is smaller than, albeit comparable with, the energy density of a free gluon plasma, due to the fact that gluon-gluon interactions are not negligible—times a (1x)(1-x) factor that accounts for the fraction that is in the deconfined phase. The evolution of the temperature of the dark sector in the primordial universe is thus determined by the competition between the heating due to the latent heat conversion into thermal energy and the adiabatic cooling induced by the Hubble expansion:
$\frac{dT}{dt}\simeq\frac{3R^{2}L_{\mbox{\tiny{h}}}T}{8\epsilon R_{0}^{3}}v-HT,$ (4.3)
where $R$ denotes the radius of the bubble, $R_{0}$ is its value at bubble percolation, which can be defined by the condition $x|_{R=R_{0}}=1/2$, and we used the fact that in general $x$ is proportional to the volume of the bubble (which we take to be spherical), so that $x=\frac{R^{3}}{2R_{0}^{3}}$, the bubble wall velocity is defined as $v=\frac{dR}{dt}$, and, following the literature, we approximated $\frac{d\epsilon}{dT}$ as $\frac{4\epsilon}{T}$, under the assumption that the $\frac{\epsilon}{T^{4}}$ ratio has only a relatively mild dependence on the temperature. Then, the condition that the temperature remains constant as a function of time leads to
$v\simeq\frac{8\epsilon R_{0}^{3}H}{3R^{2}L_{\mbox{\tiny{h}}}}.$ (4.4)
Another parameter on which the predictions for the gravitational-wave spectrum depend is the transition-strength parameter $\alpha$: it can be defined as the ratio of the trace-anomaly difference between the two phases to the thermal energy density, evaluated at the nucleation temperature (that is, at the temperature realized when the volume-averaged bubble nucleation rate reaches its maximum) [258]. It should be noted, however, that the nucleation of bubbles is a dynamical process, and the determination of the associated rate is highly non-trivial (for a detailed discussion see, for instance, refs. [259, 260, 261, 262, 263, 264, 265, 266], which describe the general aspects of this problem as well as some specific details relevant for a possible electroweak phase transition). Also in this case, our estimate for $\alpha$, that is discussed below, cannot be regarded as a fully first-principle prediction; as a consequence, the estimate for the spectrum of gravitational waves should be considered as a qualitative (or, at best, semi-quantitative) estimate, rather than a hard quantitative prediction.
Finally, a third parameter entering the determination of the gravitational-wave spectrum is the interface tension $\sigma_{\mbox{\tiny{cd}}}$ between the confining and deconfined phases. In this case, the status of the estimate for this quantity is somehow better, given that a dedicated lattice determination of $\sigma_{\mbox{\tiny{cd}}}/T_{\mbox{\tiny{c}}}^{3}$ for the $\operatorname{\mathrm{Sp}}(2)$ theory was recently published in ref. [153]. Even though the final result reported therein was obtained from simulations at fixed $N_{\tau}=4$ (and thus may be significantly different from the continuum-extrapolated result), extending the calculation that in ref. [153] was carried out for a single, relatively coarse lattice spacing—which deserved a publication on its own—to several finer lattices would require an amount of computational resources beyond the scope of the present work.
In view of these caveats, the results presented in this section should be considered as benchmark examples, rather than as genuinely first-principle predictions.
The quantity of interest to characterize the spectrum of gravitational waves is the power spectrum, defined as the fraction of the total energy density per interval of the logarithm of the frequency and denoted as $h^{2}\Omega_{\rm GW}$. Note that, for gravitational waves of cosmological origin, the minimal frequency is simply given by the Hubble rate $H$ at the time of their production. Due to the expansion of the universe, the frequency $f$ of a gravitational wave produced at time $t$ gets redshifted by the ratio of the Robertson–Walker scale factor (denoted by $a$) at time $t$ to the scale factor at the present time $t_{0}$, so that the observed frequency reads $fa(t)/a(t_{0})$. From this, one can derive an approximate estimate of the frequencies of interest for the visible sector: in the case of a first-order phase transition at the electroweak scale, of the order of $10^{2}$ GeV, taking place at a time $t\simeq 10^{-11}$ s after the Big Bang, the characteristic frequencies of the gravitational waves would be in the ballpark of $10^{-4}$ Hz. For a first-order phase transition at the scale characteristic of the strong interaction, about $10^{2}$ MeV, and taking place at the end of the so-called “quark epoch”, i.e., at $t\simeq 10^{-6}$ s, the resulting observable frequencies today would be even lower, in the range of $10^{-7}$ Hz.
@@
The generation of gravitational waves can be described by the (linearized) Einstein equation for tensor perturbations in a Friedmann background; following the notations and conventions of ref. [267], and writing the metric as $ds^{2}=a^{2}(t)\left(-dt^{2}+d{\mathbf{x}}^{2}\right)$, one has
@@
where dots denote derivatives with respect to the conformal time, $G$ is the gravitational constant, $\rho$ is the background cosmological energy density, $\Pi_{\pm}$ are the tensor helicity modes of the (dimensionless) anisotropic stress sourcing the gravitational waves, and $h_{\pm}$ represent the helicity modes of the gravitational waves.
The production of gravitational waves can take place through different mechanisms (see, for example, ref. [240, section 8] for a review): the simplest of these is the collision of true-vacuum bubbles [268, 269, 270, 271, 272, 273]. Bubbles expanding in a thermal medium, however, are subject to friction: the latent heat released in the phase transition pushes the bubbles to expand, but their interaction with the plasma constituents opposes this expansion [246]. When the two forces are balanced, the bubble walls slow down to an approximately constant speed, and the energy transferred to the plasma generates sound waves, which can be another source of gravitational waves [250, 274, 275, 276, 277, 278, 279]. In addition, the bubble-merging process can also induce hydrodynamic turbulence, which is another potential source of gravitational waves [280, 281, 282, 283, 284, 285, 286, 287].
After production, the propagation of gravitational waves is essentially free, with the phenomenologically very interesting implication that their experimental detection could provide a unique probe of the early universe, given that primordial gravitational waves decouple much earlier than photons or neutrinos. More precisely, the frequency of gravitational waves produced at the phase transition decreases as the inverse of the Robertson–Walker scale factor, while their fraction of the critical energy density decreases as $a^{-4}$. Following the discussion in ref. [38], one can then derive that entropy conservation implies that
@@
where $T_{\mbox{\tiny{CMB}}}=2.34864(49)\cdot 10^{-13}$ GeV denotes the temperature of the cosmic microwave background, while $g_{s}$ is the effective number of relativistic degrees of freedom, the $\star$ subscript refers to the time of production, while the 0 subscript refers to the time of observation.
Gravitational waves from bubble collisions. Gravitational waves produced by bubble collisions are often modelled in the thin-wall and envelope approximations [268, 37, 269, 270] (see also ref. [288]); the power spectrum can be expressed as [272, 289, 267]
$\Omega_{\mbox{\tiny{gw}}}^{\mathrm{(bc)}}h^{2}\simeq\frac{2}{3\pi}h^{2}\Omega_{\mbox{\tiny{r}}0}\Theta^{\mathrm{(bc)}}{\cal H}_{\star}^{2}t_{\mbox{\tiny{pt}}}^{2}\Omega_{S\star}^{2}v^{3}\frac{(kt_{\mbox{\tiny{pt}}})^{3}}{1+(kt_{\mbox{\tiny{pt}}})^{4}},$ (4.7)
[…]
@@
[…]
Following ref. [38], the peak frequency of the power spectrum associated with gravitational waves produced through bubble collisions in our model (for a confining transition at a temperature of the order of $500$ GeV) can be estimated to be around $1.8\cdot 10^{-4}$ Hz. Using the value of the radiation density $h^{2}\Omega_{\mbox{\tiny{r}}0}=2.47\cdot 10^{-5}$ [295], one obtains the contribution to the gravitational-wave power spectrum displayed in fig. 10, alongside those from sound waves and from hydrodynamic turbulence, that we address next.
Gravitational waves from sound waves. To discuss the spectrum of gravitational waves produced from sound waves in our $\operatorname{\mathrm{Sp}}(2)$ theory, one can follow the analysis presented in ref. [96] for a dark $\operatorname{\mathrm{SU}}(3)$ Yang–Mills model, which, in turn, relies on the formalism summarized in ref. [39] (see also refs. [275, 296, 297]).
The bubble generation and growth process involves two characteristic temperatures: the nucleation temperature and the percolation temperature. The former represents the temperature at which, on average, one bubble per Hubble time forms within a Hubble volume, while the latter can be defined as the temperature at which the probability [^7] Fixing the precise numerical value of this probability is somewhat arbitrary, and different choices are possible [298, 299, 300], but this ambiguity does not have a large impact on the estimate of the percolation temperature. of the system being in the confining vacuum (corresponding to the disappearance of supercooled regions of deconfined vacuum, that is, to the end of the phase transition) becomes comparable to one. As will be discussed below, it turns out that both the nucleation temperature and the percolation temperature are close to $T_{\mbox{\tiny{c}}}$; a heuristic argument suggesting this is based on the observation that the nucleation rate is exponentially suppressed when the temperature is decreased [301].
Other important parameters relevant for the gravitational-wave spectrum include the inverse of the product of the phase-transition duration times the Hubble parameter at the time when the gravitational waves are produced, $1/(t_{\mbox{\tiny{pt}}}H_{\star})$, as well as the ratio of the vacuum energy density released in the confining transition to the energy density of the deconfined plasma, denoted by $\alpha$. Note that the $1/(t_{\mbox{\tiny{pt}}}H_{\star})$ ratio can be estimated evaluating $\frac{d\ln\Gamma}{d\ln T}$ at the percolation temperature, where $\Gamma$ is related to the Euclidean action associated with a spherically symmetric bubble configuration at finite temperature; the larger the value of $t_{\mbox{\tiny{pt}}}H_{\star}$, the stronger the phase transition and the gravitational-wave signal. The $\alpha$ parameter, on the other hand, encodes the fraction of energy released during the phase transition that can be converted into fluid motion in the plasma, and can also be interpreted as a parameter describing the transition strength. It can be defined as
$\alpha=\frac{4}{3}\cdot\frac{\theta_{+}-\theta_{-}}{w_{+}},$ (4.9)
where $\theta=\Delta/4$, while $w=\epsilon+p$ denotes the enthalpy density, and the ++ and - subscripts respectively indicate that the quantity is evaluated just in front or just behind the bubble wall. In our setup, and under the assumption of small supercooling, one can take the values in the neighborhood of the transition temperature, in the deconfined (high-temperature) or in the confining (low-temperature) phase as proxies for these quantities. Note that, using the continuity of the pressure, eq. (4.9) can be rewritten as
$\alpha=\frac{\frac{L_{\mbox{\tiny{h}}}}{T_{\mbox{\tiny{c}}}^{4}}}{3\left(\frac{s}{T^{3}}\right)_{+}}.$ (4.10)
The results of our lattice simulations indicate that $\alpha\simeq 0.62(4)$ for the $\operatorname{\mathrm{Sp}}(2)$ Yang–Mills theory. Even though the error budget of this quantity involves systematic uncertainties that are non-trivial to quantify, taken at face value, this value would correspond to a transition that, in the terminology of ref. [302], could be described as “intermediate” to “strong”. For comparison, the value of the same parameter for the $\operatorname{\mathrm{SU}}(3)$ gauge theory is approximately $0.343$ [96].
Finally, another parameter of relevance for the study of gravitational waves produced through sound waves generated after bubble collisions is the speed of propagation of the bubble wall, evaluated in the rest frame of the fluid, at a sufficiently large distance from the bubble [303], that we denoted as $v$. For bubbles expanding in a thermal medium, such speed is expected to be limited by frictional forces and to be associated with non-trivial effects [304, 305, 306, 307, 308, 309, 310, 311, 312, 290, 255, 313, 314, 315].
@@
The power spectrum of gravitational waves produced through sound waves can be written as
@@
$h^{2}\Omega^{\mathrm{(sw)}}_{\mathrm{gw}}=2.061h^{2}F_{\mathrm{gw},0}K^{2}\left(H_{\mbox{\tiny{n}}}R_{\star}\right)\tilde{\Omega}_{\mathrm{gw}}C(f)\Upsilon(y)\Sigma,$ (4.11)
where $F_{\mathrm{gw},0}$ is the attenuation factor
$F_{\mathrm{gw},0}=3.57(5)\cdot 10^{-5}\cdot\left(\frac{100}{g_{\star,s}}\right)^{\frac{1}{3}},$ (4.12)
where $K$ is the kinetic-energy fraction of the plasma around the expanding bubbles, which can be defined from the ratio of the mean value of $w\gamma_{\mbox{\tiny{f}}}^{2}v_{\mbox{\tiny{f}}}^{2}$ (with $v_{\mbox{\tiny{f}}}$ the spatial velocity of the fluid, and $\gamma_{\mbox{\tiny{f}}}$ the associated relativistic factor) over the mean energy density of the fluid, and can be estimated from the hydrodynamics of expanding spherical bubbles [270, 274]. $K$ is computed as a function of the $H_{\mbox{\tiny{n}}}R_{\star}$ product [316], where $H_{\mbox{\tiny{n}}}$ represents the Hubble parameter at the bubble nucleation time, while $R_{\star}$ is the mean bubble separation, for which we use $R_{\star}=2\sqrt[3]{\pi}vt_{\mbox{\tiny{pt}}}$ [317]. For the $\tilde{\Omega}_{\mathrm{gw}}$ constant, we use the numerical estimate $\tilde{\Omega}_{\mathrm{gw}}=10^{-2}$ from ref. [278], from which we also take the form of the spectral shape function
$C(f)=\left(\frac{f}{f_{\mathrm{p},0}}\right)^{3}\left[\frac{7}{4+3\left(\frac{f}{f_{\mathrm{p},0}}\right)^{2}}\right]^{\frac{7}{2}},$ (4.13)
with the peak frequency
$f_{\mathrm{p},0}\simeq 26.2\left(\frac{1}{H_{\mbox{\tiny{n}}}R_{\star}}\right)\left(\frac{T_{\mbox{\tiny{n}}}}{100~\text{GeV}}\right)\left(\frac{g_{\star,s}}{100}\right)^{\frac{1}{6}}10^{-6}~\text{Hz}.$ (4.14)
The $\Upsilon(y)$ term in eq. (4.11) is related to the finiteness of the sound-wave lifetime [318] and is defined as
$\Upsilon(y)=1-\frac{1}{\sqrt{1+2y}};$ (4.15)
the argument of this function is the sound-wave lifetime in units of the Hubble time, $y=H_{\mbox{\tiny{n}}}R_{\star}/\sqrt{K}$; we take this quantity to be of the order of $10^{-2}$. The last factor on the right-hand side of eq. (4.11), $\Sigma$, accounts for attenuation effects caused by reheating [302]: it depends on the wall velocity and $\alpha$ and on $v$, and, in the most unfavorable cases, can take values as small as $10^{-3}$. For the relatively large value of $\alpha$ for our model, assuming intermediate values of $v\simeq 0.5$, we take $\Sigma\simeq 0.3$ as a crude estimate of this parameter.
Gravitational waves from turbulence. Finally, the contribution to the spectrum of gravitational waves due to hydrodynamic turbulence can be estimated as [286, 38]
$h^{2}\Omega^{\mathrm{(ht)}}_{\rm GW}$ $\simeq\frac{8}{\pi^{6}}h^{2}\Omega_{r0}\Theta^{\mathrm{(ht)}}{\cal H}_{\star}t_{\mbox{\tiny{pt}}}\Omega_{S\star}^{3/2}v^{4}\,\frac{(kt_{\mbox{\tiny{pt}}})^{3}}{\left(1+\frac{4k}{{\cal H}_{\star}}\right)\left(1+\frac{vkt_{\mbox{\tiny{pt}}}}{\pi^{2}}\right)^{11/3}},$ (4.16)
where, as in eq. (4.7), we explicitly included an effective suppression factor, set to $\Theta^{\mathrm{(ht)}}=10^{-2}$.
The gravitational-wave power spectra associated with bubble collisions, with sound waves, and with hydrodynamic turbulence are shown in fig. 10, where they are compared with the expected peak-integrated sensitivities of the LISA, DECIGO, and BBO observatories, as estimated in ref. [319]. The plot shows that the dominant contribution is expected to be the one resulting from sound waves: for the model under consideration (with the chosen parameters), this contribution to the gravitational-wave background peaks in the mHz frequency, and is predicted to be experimentally detectable. Conversely, the contributions from bubble collisions and from turbulence are expected to be subleading: they may evade detection by LISA, but could be within reach of DECIGO and BBO.
Figure 10: Power spectrum of gravitational waves produced at the first-order confining phase transition of the $\operatorname{\mathrm{Sp}}(2)$ model. The dashed, solid, and dotted concave black curves respectively represent the contributions from gravitational waves produced through bubble collisions, sound waves, and hydrodynamic turbulence, as discussed in the text, and using the estimate $\alpha\simeq 0.62(4)$ for the transition-strength parameter. In turn, the dashed red, dotted green, and dash-dotted purple curves denote the expected peak-integrated LISA, DECIGO, and BBO sensitivities [319].
@@
Before continuing the discussion of the confining phase of our model, it is appropriate to make some remarks about the parameters involved in the tentative estimate of the gravitational-wave power spectra produced at the confining transition of the $\operatorname{\mathrm{Sp}}(2)$ model. In general, the non-perturbative determination of quantities related to time-dependent phenomena and/or beyond thermodynamic equilibrium from lattice simulations is possible, but it requires techniques beyond those used for the equation of state. While Monte Carlo simulations of real-time dynamics would be hampered by a numerical sign problem, one can resort to an alternative method, focused on the study of a low-energy effective field theory describing the long-wavelength modes only [^8] A similar approach was proposed in ref. [320] and later implemented in refs. [321, 322, 323] to study the transverse-momentum broadening associated with the jet quenching phenomenon in the QCD plasma. through Langevin equations [324]: this approach was used in refs. [266, 293] to study the bubble nucleation rate (and the duration of the phase transition, which can be obtained from the latter via eq. (4.8)) at the electroweak phase transition. As discussed in ref. [38], the value of the gravitational-wave frequency corresponding to the maximum in signal intensity observable today depends on $t_{\mbox{\tiny{pt}}}$, as does the value of the signal-intensity maximum itself.